OpenStudy (anonymous):

solve the system x+6y=2 5x+4y=36 answers may be (8,-1) (2,-1) (-8,-1) (0,-8)

OpenStudy (amistre64):

-5x-30y =-10 5x +4y = 36 ------------- -26y = 26 ; y = -1

OpenStudy (anonymous):

you are too fast for me

OpenStudy (amistre64):

x = 8 then, maybe dbl chk

OpenStudy (amistre64):

x+6y=2 ; that good 5x+4y=36 40 -4 = 36

OpenStudy (amistre64):

helps to have parkinsons; then the fingers tend to type on their own ;)